# Are we there yet? A brief survey of Music Emotion Prediction Datasets, Models and Outstanding Challenges

### v1

@@ 4 Models and approaches
Some of the earliest attempts at music emotion prediction involved rule-based approaches and hierarchical frameworks. For instance, Feng et al. [67] used Computational Media Aesthetics (CMA) to analyze tempo and articulation, mapping them into four mood categories: happiness, anger, sadness, and fear. They achieved a total precision of 67% and a total recall of 66%. Lu et al. [68] developed a hierarchical framework for automatically extracting music emotion from acoustic data. They employed music intensity to represent the energy dimension of Thayer’s model while using timbre and rhythm to capture the stress dimension. They achieved an average accuracy of mood detection of up to 86.3%. These early models laid the groundwork for subsequent advancements in MER, paving the way for the adoption of more sophisticated techniques, including deep learning approaches.
In the last few years, various deep-learning MER models have been developed. Many of these leverage common techniques for time-series data, including Recursive Neural Networks (RNNs) such as Long-Short Time Memory networks (LSTMs) [69, 42], and more recently, Transformer architectures [70, 46]. Trying to identify thé current state-of-the-art MER model is a near-impossible task, due to reasons discussed in the previous section, for instance, the difference in labels across datasets. Looking at a recent competition, the 2021 Emotion and Theme Recognition in Music Using MTG-Jamendo dataset, we see that the best-performing model [66] reaches an PR-AUC-macro of 0.150872. This model uses a Convolutional neural networks (CNN) architecture.
@@
Other input modalities may include text (lyrics), or video. In the case of the former, some models extract the sentiment from the lyrics using Natural Language Processing (NLP) tools [77], or they use the entire lyrics with an embedding model [70]. These features are then combined with audio-based features or analyzed independently to predict the emotional content of music. Including the lyrics does not always improve the sentiment prediction, particularly in predicting arousal [77], however, [31] did manage to increase the performance of a MER model by using a word2vec [78] embedding trained on 1.6 million lyrics. Finally, models that include video modalities typically use pretrained networks to capture image and video features, and thus improve model performance. For instance, [45] uses ResNet-50 [79] and FlowNetS [80], respectively.
@@
5 Challenges and future directions
Whereas the first publications on music and emotion surfaced in the 1930s [50], we have not yet reached the stage where models can reach human-like performance (e.g. $>$ 95% accuracy). The field of MER still faces several challenges and various opportunities for future exploration remain, ranging from overcoming data limitations to the integration of emerging technologies. Understanding and addressing these challenges is crucial for advancing the field and unlocking its full potential.
@@ 5 Challenges and future directions
Noisy labels When creating datasets, an additional challenge arises: it can be hard to identify the emotion perceived from music, especially when working with valence and arousal models. In fact, Russel’s model [51] included a third dimension: dominance. This dimension is typically omitted because of the ambiguity in annotation. In general, categorical models, although less precise, are often easier to annotate [59]. This personal variability and ambiguity in emotion labels introduces noise in dataset labels, causing low inter-rater reliability, and making it challenging to train accurate and reliable MER models.
@@
Benchmarking The ImageNet competition has been instrumental in establishing a clear performance benchmark among computer vision models [93]. While there have been similar competitions for music emotion prediction (see Section 3), none of these competitions ran in the last three years, indicating the lack of a current benchmark for MER systems. To establish benchmarks on individual datasets, we have to revert to individual model papers, such as [28, 23]. It is unfortunately not always clear which train/test split these systems use, making direct comparisons hard. One way to facilitate an easy overview would be to leverage Leaderboard features on popular websites such as Papers With Code [^7] https://paperswithcode.com/sota and HuggingFace [^8] https://huggingface.co/docs/competitions/en/leaderboard.
@@
MIDI The forgotten format in MER. Emotion originates from many different aspects of the music, including the tonal tension [7], instrumentation and timbre [95], production quality [96], expressiveness of the performance [97], harmony [98]. The symbolic MIDI format only captures parts of these [12], which may explain the lack of models for predicting emotion from MIDI. Another reason may simply be the lack of emotion-annotated datasets (three in total). Even though MIDI is an abstraction of music, it is still a widely used format by music producers and performers and warrants its own models for emotion prediction. Even more, such datasets may enable generative music systems (which are typically trained on MIDI) to be controlled by emotion [6, 99].
@@
In sum, the challenges mentioned above provide direct opportunities and future directions to further advance the exciting field of MER.
6 Conclusion
MER is a promising field with various practical applications. With the rise of large-language models, we have seen impressive performance in various tasks. The field of MER, however, still seems to be lagging. Given the importance of large training datasets to facilitate the training of LLMs, we provide a comprehensive overview and discussion of the existing datasets for MER.
@@ 6 Conclusion
In closing, this survey serves as a valuable resource for the MER community, offering insights into the current state-of-the-art, as well as a discussion of challenges and inspiration for future directions.
7 Acknowledgements
This work has received SEED funding from SUTD TL under grant number RTDS S 22 14 04 1.

### v2

@@ IV Models and approaches
Some of the earliest attempts at music emotion prediction involved rule-based approaches and hierarchical frameworks. For instance, Feng et al. [111] used Computational Media Aesthetics (CMA) to analyze tempo and articulation, mapping them into four mood categories: happiness, anger, sadness, and fear. They achieved a total precision of 67% and a total recall of 66%. Lu et al. [112] developed a hierarchical framework for automatically extracting music emotion from acoustic data. They employed music intensity to represent the energy dimension of Thayer’s model while using timbre and rhythm to capture the stress dimension. They achieved an average accuracy of mood detection of up to 86.3%. These early models laid the groundwork for subsequent advancements in music emotion recognition, paving the way for the adoption of more sophisticated techniques, including deep learning approaches.
Figure 1: Number of Models that use Datasets in Music Emotion Recognition Models since 2020.
In recent years, various deep-learning MER models have been developed, employing techniques such as Convolutional Neural Networks (CNNs) [77, 79, 81, 82, 83, 72, 91, 84, 85, 86, 97, 104, 50, 100, 107, 110, 109], Recursive Neural Networks (RNNs) such as Gated Recurrent Units (GRUs) [100] and Long-Short Term Memory networks (LSTMs) [77, 78, 79, 80, 84, 85, 87, 97, 46, 50, 106], and more recently, Transformer architectures [88, 105, 50, 102]. Additionally, generative models such as Generative Adversarial Networks (GANs) have been applied. For instance, Huang et al. [90] developed a GAN-based model for emotion recognition using audio inputs under IoT environments.
On the Jamendo dataset, several models have demonstrated noteworthy performance. For example, Mayerl et al. [73] employed a CNN-based (ResNet18) model using clustering-based ensembles, achieving a PR-AUC-macro score of 0.1087 and a ROC-AUC-macro score of 0.7047. Tan et al. [74], using a semi-supervised learning method and a CRNN architecture, reported higher scores of PR-AUC-macro 0.1357 and ROC-AUC-macro 0.7687. Pham et al. [86] achieved similar results with an EfficientNet-based CNN model, utilizing a co-teaching strategy to manage noisy data, resulting in PR-AUC-macro 0.1435 and ROC-AUC-macro 0.7599. Bour et al. [75] leveraged frequency-dependent convolutions in a CNN model to achieve the best performance with PR-AUC-macro 0.1509 and ROC-AUC-macro 0.7748, the highest score at MediaEval 2021.
Notably, hybrid architectures combining multiple deep learning techniques, such as CNN-LSTM models, have shown promising results. For instance, Chen and Li [79] employed CNN-LSTM models for emotion recognition, outperforming standalone CNN and LSTM models. They achieved the highest accuracy of 68.1% for audio classification and 74.2% for lyric classification on the Last.fm dataset. Transformer-based models have also recently gained traction, such as those by Suresh et al. [105] and Agrawal et al. [88], which use Transformer architectures to improve MER performance. For example, the Transformer-based model proposed by Suresh et al. [105] outperformed other models, such as CNN and Bi-GRU, achieving the highest accuracy of 77.94% on the MoodyLyrics dataset.
Other models have employed more traditional machine learning techniques alongside deep learning approaches. Medina et al. [98] focused on classifying emotions from audio using SVM, Random Forest, and Multi-Layer Perceptron (MLP) models, achieving an F-score of 0.73 and 0.69 for predicting valence and arousal values, respectively, on the MediaEval dataset. Sharma et al. [76] combined machine learning algorithms such as Support Vector Machines (SVM) and Naive Bayes (NB) to predict emotions using both audio and lyric features. They achieved an accuracy of 63% on the PMEmo dataset. Similarly, Griffiths et al. [92] developed a multi-genre MER model using linear regressors. They achieved a $R^{2}$ score of 0.776 and 0.85 for predicting valence and arousal values, respectively, on the self-built dataset. Meanwhile, Xia et al. [93] utilized clustering algorithms alongside machine learning techniques like SVM and K-Nearest Neighbors (KNN) for emotion recognition, and their hybrid model which combined all the machine learning techniques achieved an accuracy of 85% on the Solymani et al.’s dataset.
@@
Recognizing the importance of induced emotions, recent research has focused on models that leverage physiological data to predict music-induced emotions. For instance, [44] constructed the HKU956 dataset with aligned peripheral physiological signals (i.e., heart rate, skin conductance, blood volume pulse, skin temperature) and self-reported emotion from 30 participants. The study revealed that physiological features significantly contribute to valence classification and that multimodal classifiers outperform single-modality ones. [89] uses PMEmo as induced emotion recognition dataset. This study merges audioLIME, a source-separation-based explainable model, with mid-level perceptual features to form an intuitive connection between input audio and emotion predictions, providing insights into model predictions. [103] also uses PMEmo as induced emotion recognition dataset and introduces a novel method named Modularized Composite Attention Network (MCAN). This method enhances feature extraction and employs attention mechanisms to improve the stability and accuracy of emotion prediction models.
@@
Other input modalities may include text (lyrics), or video. In the case of the former, some models extract the sentiment from the lyrics using Natural Language Processing (NLP) tools [99, 96], or they use the entire lyrics with an embedding model [105, 101]. These features are then combined with audio-based features or analyzed independently to predict the emotional content of music. Including the lyrics does not always improve the sentiment prediction, particularly in predicting arousal [99]; however, [108] did manage to increase the performance of a MER model by using embeddings, such as word2vec [117] or stacked ensemble models that integrate both audio and lyrics. Finally, models that include video modalities typically use pretrained networks to capture image and video features, and thus improve model performance. For instance, [49] uses ResNet-50 [118] and FlowNetS [119], respectively.
Identifying the current state-of-the-art Music Emotion Recognition (MER) model is challenging due to factors such as differences in datasets and performance metrics, as highlighted in the previous section. One standard evaluation is provided by the 2021 Emotion and Theme Recognition in Music competition, which uses the MTG-Jamendo dataset. The best performing model achieved a PR-AUC-macro of 0.150872 with a convolutional neural network (CNN) architecture [75].
Figure 1 shows which datasets have most been used in Music Emotion Recognition models since 2020, as presented in Table II. As shown, self-built datasets, which are created by the authors themselves for specific research purposes, are the most commonly used, appearing in 11 instances. These datasets may contain licenced data and are not available to the research community. Other frequently utilized datasets include DEAM (8 occurrences) and MTG-Jamendo (6 occurrences), indicating their popularity in MER research. Some datasets, such as PMEmo and Solymani et al.’s dataset, appear less frequently but still contribute valuable insights.
@@
V Challenges and future directions
Whereas the first publications on music and emotion surfaced in the 1930s [58], current MER models still struggle to match human performance in emotion recognition. The field of MER still faces several challenges and various opportunities for future exploration remain, ranging from overcoming data limitations to the integration of emerging technologies. Understanding and addressing these challenges is crucial for advancing the field and unlocking its full potential.
@@ V Challenges and future directions
Noisy labels When creating datasets, an additional challenge arises: it can be hard to identify the emotion perceived from music, especially when working with valence and arousal models. In fact, Russell’s model [59] included a third dimension: dominance. This dimension is typically omitted because of the ambiguity in annotation. In general, categorical models, although less precise, are often easier to annotate [68]. This personal variability and ambiguity in emotion labels introduces noise in dataset labels, causing low inter-rater reliability, and making it challenging to train accurate and reliable MER models. In addition, some emotion representations are prone to more noise. For instance, the MTG-Jamendo dataset [36] offers a large number of freely assigned tags, making it suitable for exploring a wide range of emotions, but the lack of a controlled annotation process can lead to noisy data, as there may be synonyms of emotion terms.
@@
Benchmarking The ImageNet competition has been instrumental in establishing a clear performance benchmark among computer vision models [132]. While there have been similar competitions for music emotion prediction (see Section III), none of these competitions ran in the last three years, indicating the lack of a current benchmark for MER systems. To establish benchmarks on individual datasets, we have to revert to individual model papers, such as [32, 24]. It is unfortunately not always clear which train/test split these systems use, making direct comparisons hard. One way to facilitate an easy overview would be to leverage Leaderboard features on popular websites such as Papers With Code [^7] https://paperswithcode.com/sota and HuggingFace [^8] https://huggingface.co/docs/competitions/en/leaderboard.
@@
MIDI The forgotten format in music emotion recognition. Emotion originates from many different aspects of the music, including the tonal tension [134], instrumentation and timbre [135], production quality [136], expressiveness of the performance [137], harmony [138]. The symbolic MIDI format only captures parts of these [12], which may explain the lack of models for predicting emotion from MIDI. Another reason may simply be the lack of emotion-annotated datasets (three in total). Even though MIDI is an abstraction of music, it is still a widely used format by music producers and performers and warrants its own models for emotion prediction. Even more, such datasets may enable generative music systems (which are typically trained on MIDI) to be controlled by emotion [6, 139].
@@
In sum, the challenges mentioned above provide direct opportunities and future directions to further advance the exciting field of music emotion recognition.
VI Conclusion
Music emotion recognition is a promising field with various practical applications. With the rise of large-language models, we have seen impressive performance in various tasks. The field of music emotion recognition, however, still seems to be lagging. Given the importance of large training datasets to facilitate the training of LLMs, we provide a comprehensive overview and discussion of the existing datasets for music emotion recognition.
@@ VI Conclusion
In closing, this survey serves as a valuable resource for the MER community, by offering insights into the current state-of-the-art, as well as a discussion of challenges and inspiration for future directions.
VII Acknowledgements
This work has received SEED funding from SUTD TL under grant number RTDS S 22 14 04 1 and SUTD SKI 2021_04_06.